OpenStudy (dls):

Integrate sqrt(x^2+4) (x^2+1)

OpenStudy (dls):

\[\Huge \sqrt(x^2+4)(x+1)\] I am thinking of subbing x=1/t not sure didn't get me anywhere

ganeshie8 (ganeshie8):

wolfram gave that

OpenStudy (dls):

i saw the WF solution,please give me a better solution :/